Suppose that you and a friend are playing cards and you decide to make a friendly wager. The bet is that you will draw two cards without replacement from a standard deck. If both cards are hearts, your friend will pay you $25.00. Otherwise, you have to pay your friend $5.00. What is the expected value of your bet?
Answer #1

Answer:

Drawing two hearts (without replacement) from a standard deck ,Probability = 13 / 52 * 12 / 51

= 0.0588

Expected value of the bet

= (0.0588*25)+(1-0.0588)(-5)

= 1.47+[0.9412*(-5)]

= 1.47-4.706

= -3.236

The Expected loss = 3.236 dollars.